OpenStudy (anonymous):

Write a quadratic equation in the variable x having the give numbers as solution. Type the equation in standard form, ax^2+ bx+c=0 solution:10, only solution

OpenStudy (llort):

pardon? was this cut off?

OpenStudy (dumbcow):

y = (x-10)^2 --> x^2 -20x + 100 = 0

OpenStudy (llort):

now I see it, there is only one intersect, (x+10)(x+10) X^2+20X+100=0

OpenStudy (anonymous):

dumbcow got it. If the quadratic only had 10 as a solution, when factored it would have to be: \[(x-10)(x-10) = (x-10)^{2} = x^{2}-20x+100\]

OpenStudy (llort):

yep, the methane producing cud-chewer got it again...

OpenStudy (anonymous):

same but 3, only solution please

OpenStudy (dumbcow):

just plug in 3 where 10 was before :)

OpenStudy (llort):

(x+3)(x+3) go from there...

OpenStudy (anonymous):

its (x - 3)(x - 3)

OpenStudy (llort):

give a man a fish, he'll eat for a day; teach a man to fish, he'll eat for a lifetime.

OpenStudy (anonymous):

x^2-30x+1000

OpenStudy (anonymous):

is that correct

OpenStudy (llort):

remember FOIL First Outside Inside Last

OpenStudy (anonymous):

(x-3)(x-3)=(x-3)^2=(x-3)^2-6x+9 is that correct

OpenStudy (llort):

First: x*x=x^2 Outside: x*-3=-3x Inside: -3*x=-3x Last: -3*-3=9 now you have: x^2-3x-3x+9 Which can simplify to: x^2-6x+9

OpenStudy (llort):

do you get it?